Circular No. 5839

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERSEID METEORS Following the impressive Perseid activity in 1991 and 1992, with hourly

rates of a few hundred for ∼ 1 hour near the time t_{0} when the earth was crossing the orbit plane of P/Swift-Tuttle (IAUC 5330, 5386), and the subsequent recovery of the comet (IAUC 5620), there has been widespread speculation on the possibility of even a storm of Perseids (e.g., Rao 1993, Sky Tel. 86, No. 2, 43) around t_{0} = 1993 Aug. 12.05 UT, some 7.5 months after the comet was at its descending node. Although records of both the comet (IAUC 5670) and the Perseids have been traced back for 2000 years, there is no record of a Perseid shower prior to the nineteenth century within several years of a passage of the comet through this node. There was impressive Perseid activity during 1861-1863, a dynamical difference then being that the distance between the node and the earth’s orbit, 0.005 AU, was significantly less than the 0.01-0.03 AU at the comet’s previous nodal passages (Marsden et al. 1993, Icarus, submitted). It is the present even smaller distance, 0.0009 AU, coupled with the circumstantial evidence that meteor storms tend to follow their parent comets (cf. the Leonids; e.g., Yeomans 1981, Icarus 47, 492), that has led to the current interest. On the assumption that the 1991 and 1992 events were produced by the ejection of meteoroids from P/Swift-Tuttle in 1862, I. P. Williams and Z. Wu, Queen Mary and Westfield College, London, calculate that the maximum activity in 1993 could be at any time between Aug. 11.92 and 12.06 UT, and that there is more likelihood of an actual Perseid storm in 1994. Ľ. Kresák, Astronomical Institute, Bratislava, indicates that any real storm in 1993 would be likely to occur close to t_{0}. In any case, comparison with 1863 (Olson and Doescher 1993, Sky Tel. 86, No. 2, 47), when t_{0} = 1863 Aug. 10.68 UT occurred some 11 months after the comet passed its node, suggests the potential for impressive Perseid activity up to fully 0.7 day after t_{0}.

R. Stachnik, NASA Headquarters, draws attention to the possibility of observing sunlight scattered from meteoritic dust in the general direction of the Perseid radiant and anti-radiant. Rao (loc. cit.) indicates that the points to observe are ∼ 10^{o} south of Algol and in Triangulum Australe. J. G. Hills and C. Snell, Los Alamos National Laboratory, remark that, if the earth experiences a meteor storm, many meteoroids would hit the moon ∼ 0.15 day earlier. They suggest that there is a non-negligible probability of impact by 1-m meteoroids, which with densities of 0.2 g cm^{−3} and impact velocity 60 km/s could produce observable flashes of duration 1 second to mag 1 on the dark portion of the last-quarter moon.

1993 August 5 Brian G. Marsden